OpenStudy (anonymous):

Jordan and Taylor went to a baseball game and bought the same number of sodas and hotdogs throughout the game, but they went to different food vendors. Jordan’s vendor charged $2 per soda, $5 per hotdog, and he spent $23. Taylor’s vendor charged $3 per soda, $4 per hotdog, and she spent $24. How many sodas did each of them buy? 1 2 3 4

OpenStudy (anonymous):

a: # of soda bought b: # of hotdogs bought \[2a+5b=23\]\[3a+4b=24\] \[6a+15b=69\]\[6a+8b=48\] \[7b=21\] \[b=3\] We want a, so but b in one of the equation. \[2a+5*3=23\]\[ 2a=8 \]\[a=4\]